# DIABLO: A 6-DoF Wheeled Bipedal Robot Composed Entirely of Direct-Drive Joints

### v2

@@ IV MODELING
As illustrated in Fig. 5, a local control frame ($o_{n}$) is positioned at the midpoint of the line connecting the two wheels. The body frame ($o_{b}$) is established relative to the local control frame. The body frame’s origin is positioned at the center of the IMU, which is installed inside the head. Specifically, the IMU is situated at the midpoint of the centerline connecting the left and right hip joints. Prior to employing the Newton-Euler method for analysis, the leg linkage structure was initially simplified to a variable-length straight rod, and the model was further reduced to a 2D plane. The pitch, yaw, and roll values are obtained from the rotation of the body frame in relation to the inertia frame.
@@
${}_{I}x_{H}$ $={}_{I}x+Lsin\theta+lcos(\beta+\theta)$ (1)
${}_{I}z_{H}$ $=Lcos\theta+lsin(\beta+\theta)$
where ${}_{I}x$ is the displacement of the $o_{n}$ system relative to the $o_{i}$ system in the x-direction; $L$ is the length of the rod; $\theta$ is the pendulum angle of the rod; $l$ is the distance from the CoM1 to the hip joint center; $\beta$ is the angle between the CoM1 and the head’s x-axis of rotation.
@@
$I_{L}\ddot{\theta}=\tau_{w}-\tau_{H}$ $+(F_{z1}\sin\theta-F_{x1}\cos\theta)L_{w}$
@@
$I_{H}(\ddot{\theta}+\ddot{\beta})=\tau_{H}-F_{z1}lcos(\theta+\beta)-F_{x1}lsin(\theta+\beta)$
@@
where $\bm{q}=[\gamma,\theta,x]^{T},\bm{u}=[\tau_{H},\tau_{w}]^{T}$, $\gamma$ is the pitch angle of the head of DIABLO in the 2D model, and the controlled state variable $\bm{q}$ must be obtained through the transformation of joint position state variables. The corresponding relationship is given by:

### v3

@@ IV MODELING
As illustrated in Fig. 5, a local control frame ($o_{n}$) is positioned at the midpoint of the line connecting the two wheels. The body frame ($o_{b}$) and simple rod frame ($o_{r}$) are established relative to the local control frame. The body frame’s origin is positioned at the center of the IMU, which is installed inside the head. Specifically, the IMU is situated at the midpoint of the centerline connecting the left and right hip joints. Prior to employing the Newton-Euler method for analysis, the leg linkage structure was initially simplified to a variable-length straight rod, and the model was further reduced to a 2D plane. The pitch, yaw, and roll values are obtained from the rotation of the body frame in relation to the inertia frame.
@@
${}_{I}x_{H}$ $={}_{I}x+Lsin\theta+lcos(\beta+\theta+\phi)$ (1)
${}_{I}z_{H}$ $=Lcos\theta+lsin(\beta+\theta+\phi)$
where ${}_{I}x$ is the displacement of the $o_{n}$ system relative to the $o_{i}$ system in the x-direction; $L$ is the length of the rod; $\theta$ is the pendulum angle of the rod; $l$ is the distance from the CoM1 to the hip joint center; $\beta$ represents the angle of rotation of the $o_{b}$ frame relative to the $o_{r}$ frame; $\phi$ is the angle between the CoM1 and the head’s x-axis of rotation($x_{b}$).
@@
$I_{L}\ddot{\theta}=-\tau_{w}-\tau_{H}$ $+(F_{z1}\sin\theta-F_{x1}\cos\theta)L_{w}$
@@
$I_{H}(\ddot{\theta}+\ddot{\beta})=\tau_{H}-F_{z1}lcos(\theta+\beta+\phi)-F_{x1}lsin(\theta+\beta+\phi)$
@@
where $\bm{q}=[\gamma,\theta,x]^{T},\bm{u}=[\tau_{H},\tau_{w}]^{T}$, $\gamma$ is the pitch angle of the head of DIABLO in the 2D model, which satisfies equation $\gamma=\beta+\theta$. The controlled state variable $\bm{q}$ must be obtained through the transformation of joint position state variables. The corresponding relationship is given by: